# Necessary and Sufficient Conditions for Proving Choice in Zermelo-Fraenkel Set Theory

### v1

@@ 3 The Theorem of Choice for Families of Well-Ordered Sets
Alternatively, we can start from the fact that there exists a non-empty set of well-orders over $A$, denoted $\mathcal{R}_{A}$, thus the following statement is true:
$\forall{A\in A_{S}}\exists{R_{A}\in\mathcal{P}(A\times A)}(R_{A}\in\mathcal{R}_{A}).$
Since $A_{S}\neq\emptyset$, by the logical property that $\forall{x\in A_{S}}(\psi(x))\implies\exists{x\in A_{S}}(\psi(x))$, we can imply the statement above remains true if $\forall A$ is replaced with $\exists A$:
$\exists{}{A\in A_{S}}\exists{R_{A}\in\mathcal{P}(A\times A)}(R_{A}\in\mathcal{R}_{A}).$
Then, since $\mathcal{P}(A\times A)\subseteq U_{1}$ and $\exists{x\in X}(\psi(x))\land{X\subseteq Y}\implies\exists{y\in Y}(\psi(y))$, and defining $R_{A}=R\cap(A\times A)$), we infer that
$\exists{A\in A_{S}}\exists{R\in U_{1}}(R_{A}\in\mathcal{R}_{A}),$
and, rearranging the quantifies,
$\exists{R\in U_{1}}\exists{A\in A_{S}}(R_{A}\in\mathcal{R}_{A}).$
Using the bijection $p$, we can map $R$ to the relevant $Q=p(R)$, and $Q_{A}=p(R_{A})$:
$\exists{Q\in U_{2}}\exists{A\in A_{S}}(p^{-1}(Q_{A})\in\mathcal{R}_{A})$
and, since $p^{-1}(Q_{A})\in\mathcal{R}_{A}$, there should exist a well-order $R_{A}$ that maps to $Q_{A}$:
$\exists{Q\in U_{2}}\exists{A\in A_{S}}\exists{R_{A}}(p^{-1}(Q_{A})=R_{A}),$
which is equivalent to
$\lnot\forall{Q\in U_{2}}\forall{A\in A_{S}}\forall{R_{A}}(p^{-1}(Q_{A})\neq R_{A}),$
and, since $A_{S}\neq\emptyset$, we can replace $\forall{A\in A_{S}}$ with $\exists{A\in A_{S}}$:
$\lnot\forall{Q\in U_{2}}\exists{A\in A_{S}}\forall{R_{A}}(p^{-1}(Q_{A})\neq R_{A}).$
Note that the statement above is exactly the negation of (16), and since this statement holds true, (16) must be false, so $Q_{S}\neq\emptyset$.
@@ 5 Necessary and Sufficient Conditions for Proving Choice in ZF without AC
Thus, it only remains to prove that $\mathcal{R}_{A}\neq\emptyset$ or, equivalently, that $\varphi_{3}(R,A,f_{c})$ is not false for every possible $R\in\mathcal{P}(A\times A)$. To do this, let’s consider two binary relations over a given $A$. First, the identity function, $id_{A}=\big\{(a,a)|a\in A\big\}$—it is clear that $id_{A}\subset A\times A$ and, since $A\neq\emptyset$, we have $id_{A}\neq\emptyset$. Second, a binary relation generated on a given set $A$ using a given function of choice $f_{c}$, $f_{A,f_{c}}=\big\{(a,b)|a=f_{c}(A)\land b\in A\big\}$—it is clear that $f_{A,f_{c}}\subset A\times A$ and, since $\exists{f_{c}}\land A\neq\emptyset$, we have $f_{A,f_{c}}\neq\emptyset$. Now, for any $A\in A_{S}$ and any function of choice $f_{c}$, there exists a union $(id_{A}\cup f_{A,f_{c}})\subset A\times A$, and it is quite evident that $\varphi_{3}(id_{A}\cup f_{A,f_{c}},A,f_{c})$ is true. Therefore, $\mathcal{R}_{A}\neq\emptyset$. ∎

### v2

@@ 3 The Theorem of Choice for Families of Well-Ordered Sets
Alternatively, we can note that (16) is equivalent to
$\forall{R^{\prime}\in U_{1}}\big[\exists{A\in A_{S}}\forall{R_{A}\in\mathcal{R}_{A}}\big(R^{\prime}_{A}\neq R_{A}\big)\big],$
or $\forall{R^{\prime}\in U_{1}}\big[\exists{A\in A_{S}}\big(R^{\prime}_{A}\notin\mathcal{R}_{A})\big]$, where $R^{\prime}_{A}=R^{\prime}\cap(A\times A)$ is a restriction of $R^{\prime}$ to $A$. Since $U_{1}$ by construction contains all possible orderings for every $A\in A_{S}$, but none of its elements belongs to $\mathcal{R}_{A}$ for some $A$, there are no orderings in $\mathcal{R}_{A}$, i.e. $\exists{A\in A_{S}}(\mathcal{R}_{A}=\emptyset)$, which contradicts to the conditions of the theorem, so our initial assumption must be false, and $Q_{S}\neq\emptyset$.
@@ 5 Necessary and Sufficient Conditions for Proving Choice in ZF without AC
Thus, it only remains to prove that $\mathcal{R}_{A}\neq\emptyset$ or, equivalently, that $\varphi_{3}(R,A,f_{c})$ is not false for every possible $R\in\mathcal{P}(A\times A)$. To do this, let’s consider two binary relations over a given $A$. First, the identity function, $id_{A}=\big\{(a,a)|a\in A\big\}$—it is clear that $id_{A}\subset A\times A$ and, since $A\neq\emptyset$, we have $id_{A}\neq\emptyset$. Second, a binary relation generated on a given set $A$ using a given function of choice $f_{c}$, $f^{-1}_{A,f_{c}}=\big\{(a,b)|a=f_{c}(A)\land b\in A\big\}$—it is clear that $f^{-1}_{A,f_{c}}\subset A\times A$ and, since $\exists{f_{c}}\land A\neq\emptyset$, we have $f^{-1}_{A,f_{c}}\neq\emptyset$. Now, for any $A\in A_{S}$ and any function of choice $f_{c}$, there exists a union $(id_{A}\cup f^{-1}_{A,f_{c}})\subset A\times A$, and it is quite evident that $\varphi_{3}(id_{A}\cup f_{A,f_{c}},A,f_{c})$ is true. Therefore, $\mathcal{R}_{A}\neq\emptyset$. ∎